On January 1, 2021, the general ledger of 3D Family Fireworks includes the following account balances: Credit Debit $ 26,100View transaction list View journal entry worksheet General Journal Credit No 1 Date Jan 02, 2021 Debit 46,100 Cash Service ReJournal entry worksheet 9 13 > The company estimates future uncollectible accounts. The company determines $4,000 of accountsJournal entry worksheet < 0 ..... @ 6 6 6 13 > Supplies at the end of January total $800. Record the adjusting entry for suppJournal entry worksheet < 0 ..... 6 6 8 9 10 13 Accrued interest revenue on notes receivable for January. Interest is expecteJournal entry worksheet < 0..... 6 0 6 9 10 11 13 > Unpaid salaries at the end of January are $34,600. Record the adjusting eView transaction list Journal entry worksheet < 0 ..... 6 0 8 9 10 11 12 13 > Record the closing entry for revenue. Note: EntJournal entry worksheet < 0..... 6 0 9 10 11 12 13 ) Record the closing entry for expenses. Note: Enter debits before credits
Answer #1

Accounts Receivable as on Jan 1 = $14700+83400-2500-81000 = $14600
Allowance for Doubtful accounts unadjusted = $3000-2500 = $500
Allowance for Doubtful accounts Adjusted = 4000 x 20% + (14600-4000)x5% = $1330

Account Titles Debit Credit
8 Bad Debt Expense $                830 =1330-500
       Allowance for Doubtful Debts $                    830
9 Supplies Expense $            2,800 =3600-800
     Supplies $                2,800
10 Interest Receivable $                  75 =15000*6%*1/12
      Interest Revenue $                      75
11 Salaries Expense $          34,600
      Salaries Payable $              34,600
12 Service Revenue $       129,500
Interest Revenue $                  75
      Retained Earnings $           129,575
13 Retained Earnings $          85,000
       Salaries Expense $              67,100
       Utilities Expense $              14,800
       Bad Debt Expense $                    300
      Supplies Expense $                2,800